OpenStudy (anonymous):

Simplify the following @ganeshie8

OpenStudy (anonymous):

\[3^{n+1}-3^{n-1}\]

OpenStudy (freckles):

well you can factor a 3^n out

OpenStudy (anonymous):

\[3^n~\times~3^1-3^n~\div~3^1\]

OpenStudy (anonymous):

Combining like terms

OpenStudy (freckles):

\[3^n(3-\frac{1}{3})\]

OpenStudy (freckles):

now just find the difference of the fractions in the ( )

OpenStudy (anonymous):

\[3^n(\frac{ 8 }{ 3 })\]\[8(3^{n-1}) ?\]

OpenStudy (freckles):

those are both right I don't know which of the ones you and your teacher find most simplified

OpenStudy (freckles):

what do you mean

OpenStudy (freckles):

we show our work I thought

OpenStudy (freckles):

are you saying the final answer is suppose to be 8(3^(n-1)) ? if so all you need is quotient rule

OpenStudy (freckles):

\[\frac{3^n}{3}8 =\frac{3^n}{3^1}8 =3^{n-1} 8\]

OpenStudy (anonymous):

Thnx @freckles

OpenStudy (freckles):

also here is another way: \[3^{n+1}-3^{n-1} \\ 3^{n-1+2}-3^{n-1} \\ 3^23^{n-1}-3^{n-1} \\ 9(3^{n-1})-1(3^{n-1}) \text{you have like terms } \\ 8(3^{n-1})\]

OpenStudy (freckles):

either way @MARC_ \[8(3^{n-1})=8 \frac{3^n}{3^1} \text{ by quotient rule } \\ 8 \frac{3^n}{3} \text{ since } 3^1=3 \\ \frac{8}{3} 3^n \]

OpenStudy (anonymous):

alright.Thnx @freckles :D